OpenStudy (anonymous):

Match each function with the expression representing its inverse function. 1. y = x - 4 x + 4 2. g(x) = -0.5x -2x 3. k(x) = x x - 4 4. p(x) = x + 4 2x 5. h(x) = 4x 0.25x 6. ƒ(x) = x/2 x

OpenStudy (anonymous):

hey i just need a little help matching these

OpenStudy (butterflydreamer):

do you know how to find an inverse function?

OpenStudy (anonymous):

no i cant remember

OpenStudy (anonymous):

its y = f(x) right?

OpenStudy (butterflydreamer):

okay basically, you start off with a function. Then you switch around the x values to y values and vice versa. Then solve for y. E.g. in your first function, we have: y = x - 4 To find the inverse, change the y to an x and the x to a y. So, x = y - 4 Now we solve for y by adding 4 to both sides. Therefore, x + 4 = y Therefore your ORIGINAL function was y = x - 4 , the INVERSE function is y = x + 4

OpenStudy (butterflydreamer):

make sense??

OpenStudy (anonymous):

oh so you pretty much switch then solve

OpenStudy (butterflydreamer):

yep :) and you do the same for all of them ^_^

OpenStudy (anonymous):

and yes thx for the example it really helps in comparison to the people that just give me the answer

OpenStudy (butterflydreamer):

no probleeem ^_^ hhahaha glad you liked the example! Do you know where to go from here?

OpenStudy (anonymous):

yes im almost done with the 3rd one. :)

OpenStudy (butterflydreamer):

oh, that's really good! ^_^ just ask if you get stuck LOL

OpenStudy (anonymous):

would you like me to write you a testimonial?

OpenStudy (butterflydreamer):

LOOOOL only if you want to :)

OpenStudy (anonymous):

cool alright im gonna get the rest of these done

OpenStudy (anonymous):

ok im finished and i got them all right on odysseyware. :)

OpenStudy (butterflydreamer):

ahh GREAT! :D